Calculate (a) the number of moles in 1.00m^3 of gas at STP, and (b) the number of liters of gas per mole.
Answer #1

(a) At STP 1.0 mol =22.4 Liters

Since 1.0 m3=1000 Liters

So the number of moles in 1.00m3 n=1000/22.4

n=44.64 moles Answer

b) Using gas equation V=nRT/P ..1

Here

At STP T=273.15 K ,P=1.013*105 Pa , R=8.314 J/mol.K ,n=1 mol

Put these value in equation 1

V=(1*8.314*273.15)/(1.013*105)

V=2241.82*10-5

V=0.0224182 m3

since 1.0 m3=1000 Liters

so

V=0.0224182*1000

V=22.4182 Liters

V=22.4 Liters in 1.00 mol